OpenStudy (anonymous):

HELP PLEASE!! Susan keeps track of the number of tickets sold for each play presented at the community theater. within how many standard deviations of the mean do all the values fall? 135, 71, 69, 80, 158, 152, 161, 96, 122, 118, 87, 85.

OpenStudy (anonymous):

First you need to find the mean (the average) Do you know how to do that ?

OpenStudy (anonymous):

yeah the mean=111.16

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

read this.... http://www.mathsisfun.com/data/standard-deviation.html

OpenStudy (anonymous):

so the mean is wrong

OpenStudy (anonymous):

135 + 71 + 69 + 80 + 158 + 152 + 161 + 96 + 122 + 118 + 87 + 85 = 1334. Now take 1334 and divide it by the total # of numbers, which is 12. 1334/12 = 111.166 rounds to 111.17 Now take each numbers difference from the mean and square it. 135 - 111.17 = 23.83 71 - 111.17 = - 40.17 69 - 111.17 = -42.17 80 - 111.17 = -31.17 158 - 111.17 = 46.83 152 - 111.17 = 40.83 161 - 111.17 = 49.83 96 - 111.17 = - 15.17 122 - 111.17 = 10.83 118 - 111.17 = 6.83 87 - 111.17 = - 24.17 85 - 111.17 = - 26.17 28.83^2 + (-40.17)^2 + (-42.17)^2 + (-31.17)^2 + 46.83^2 + 40.83^2 + 49.83^2 + (-15.17)^2 + 10.83^2 + 6.83^2 + (-24.17)^2 + (-26.17)^2 = 831.17 + 1613.63 + 1778.31 + 971.57 + 2193.05 + 1667.09 + 2483.03 + 230.13 + 117.29 + 46.65 + 584.19 + 684.87 = 13200.98. now take 13200.98 and divide it by 12 = 1100.08 (this is your variance) The standard deviation is the square root of your variance. sq rt of 1100.08 = 33.17 I am going to say the standard deviation is 33.17. If I added all of this correctly. You might want to get another opinion.

OpenStudy (anonymous):

@Opcode ....can you tell me if I did this correctly. I am not asking you to do the math, just tell me if I set it up right

OpenStudy (anonymous):

for answers I have 3, 4, 5, 2 so the answer's 3??

OpenStudy (anonymous):

oh...I must have messed up big time....I don't know....let me get somebody here to help

OpenStudy (anonymous):

@Luis_Rivera ....I must have messed up big time on this...can you help

OpenStudy (anonymous):

@Preetha ....can you help

OpenStudy (anonymous):

@jim_thompson5910 ......I think I messed up here....pretty sure I did...can you help

OpenStudy (anonymous):

@satellite73 ....I really messed up on this one....can you help fix it ?

OpenStudy (anonymous):

@saifoo.khan ....please help

OpenStudy (anonymous):

@Chooch146 ....I messed up here, can you help me fix it ?

OpenStudy (anonymous):

what are your choices?

OpenStudy (anonymous):

work looks right, answer is here http://www.wolframalpha.com/input/?i=standard+deviation+ {135%2C+71%2C+69%2C+80%2C+158%2C+152%2C+161%2C+96%2C+122%2C+118%2C+87%2C+85.}

OpenStudy (anonymous):

2,3,4,5.....that is what nicole said the answer choices were

OpenStudy (anonymous):

looks like sd is 34.3

OpenStudy (anonymous):

Its d within two standard deviations. When you calculate standard deviation for those number you get 36.35 and the mean is 116.1. all of those numbers are contained within 116.1 +or - 2(36.35)

OpenStudy (anonymous):

So the answer is 2 ?

OpenStudy (anonymous):

um yup!

OpenStudy (anonymous):

thank you......you are a miracle worker

OpenStudy (anonymous):

aw lolol thanks:P

OpenStudy (anonymous):

@nicole-0109 .....Chooch got the answer....it is 2

OpenStudy (anonymous):

thank you so much to all

OpenStudy (anonymous):

:P